A wire conductor in the shape of an “L” of square sides ? = 0.300 m carries a current I = 20.0 A. Determine the magnetic field at a point x located at the center of the “L” (at x = 0.150 m above the horizontal wire and to the right of the vertical wire).

A wire conductor in the shape of an “L” of square
Answer #1

= angle by each end of the wire at the point = tan-1((l/2)/0.150) = tan-1((0.15)/0.150) = 45

r = perpendicular distance of point from each wire = 0.15

B = magnetic field by each wire = (/4) (i/r) (Sin + Sin) = (10-7) (20/0.15) (2 Sin45) = 1.886 x 10-5 T

net magnetic field by the two wires = Bnet = 2B = 2 (1.886 x 10-5 ) = 3.77 x 10-5 T